Diffusion Morning Report 04/07/11

Ted Berg tberg at stanford.edu
Thu Apr 7 06:29:04 PDT 2011


Good Morning All,
     Looking good this morning. Epi had some SiH4 flow and Mapping 
issues. Tylan LTO had O2 flow issues. It looks like we will be starting 
the PM on 4160 this morning if FS shows up. Thermco 2 no TLC flow. Ebeam 
compressor is dead. no other issues at this time.

					
	Date: 	4/7/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*.88/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*.12/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*9.92/30* 	** 	**
Tylan 11-LTO 	*yellow* 	*O2 flow issues* 	*8.38/30* 	** 	*1*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*yellow* 	*mapping and SiH4 flow issues* 	** 	** 	*2*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 479 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110407/47190031/attachment.html>


More information about the diffusion mailing list